Assume that A and B are events in a probability space with the property that P(A) = 0.5, P(B) = 0.6, and P(A ∪ B) = 0.9.

1. Explain why A and B cannot be independent.

2. Is A favorable or unfavorable to B? (Remember that an event E is said to be favorable to F if P(F|E) > P(F); that is, if the knowledge that E occurred increases the plausibility of F.)
Answer #1

Given P (A) 0.5. P (B) 0.6 and P\left ( A\cup B \right )=0.9 .

Now

P\left ( A\cup B \right )=P\left ( A \right )+P\left ( B \right )-P\left ( A\cap B \right )\\ 0.9=0.5+0.6-P\left ( A\cap B \right )\\ P\left ( A\cap B \right )=0.2

a) If A and B are independent, thenP\left ( A\cap B \right )=P\left ( A \right )P\left ( B \right ). Here

P\left ( A\cap B \right )=0.2\neq 0.3=0.5\times 0.6=P\left ( A \right )P\left ( B \right )

Hence A and B cannot be independent.

b) We have using conditional probability P\left ( B|A \right )=P\left ( A\cap B \right )/P\left ( A \right )=0.2/0.5=0.4

Since P\left ( B|A \right )=0.4<0.6=P\left ( B \right ) , A is unfavorable to B.